OpenStudy (anonymous):

Given the exponential equation 2x = 128, what is the logarithmic form of the equation in base 10?

OpenStudy (aum):

\(\Large 2^x = 128\) Take logarithm to the base 10 on both sides. What do you get?

OpenStudy (anonymous):

umm idk how do I find that?

OpenStudy (aum):

\(\Large \log_{10}2^x = \log_{10}128\) Use the logarithm property: \(\Large \log(a^n) = n * \log(a)\)

OpenStudy (anonymous):

log base 10 of 2, all over log base 10 of 128 ?

OpenStudy (jdoe0001):

\(\large 2^x = 128?\)

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

\(\large \bf log_{\color{red}{ a}}{\color{blue}{ b}}=y\implies {\color{red}{ a}}^y={\color{blue}{ b}} \\ \quad \\ \quad \\ {\color{red}{ 2}}^x = {\color{blue}{ 128}}\implies ?\)

OpenStudy (aum):

\(\Large \log_{10}2^x = \log_{10}128\) \(\Large x\log_{10}2 = \log_{10}128\)

OpenStudy (jdoe0001):

anyhow \(\bf log_{\color{red}{ a}}{\color{blue}{ b}}=y\implies {\color{red}{ a}}^y={\color{blue}{ b}} \\ \quad \\ \quad \\ {\color{red}{ 2}}^x = {\color{blue}{ 128}}\implies log_{\color{red}{ 2}}{\color{blue}{ 128}}=x \\ \quad \\ \textit{then you can just use the change of base rule }log_ab=\cfrac{log_{\color{olive}{ c}}b}{log_{\color{olive}{ c}}a}\) for the log change of base rule, "c" can pretty much be anything, so long is the same atop and below

OpenStudy (anonymous):

log_(2)128/log_(2)2?

OpenStudy (jdoe0001):

well... you're expected to use log base10 though

OpenStudy (aum):

Are you asked to solve for x or just write the logarithmic form of the equation to the base 10?

OpenStudy (anonymous):

oh thank you haha yah so instead of 2 its ten

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Ahhhh thankyou soo much!! This was corrrect ;)